A property was purchased for

$9127.00

down and payments of

$768.00

at the end of every

year

for

9

years. Interest is

9 %

per annum compounded

quarterly

What was the purchase price of the property? How much is the cost of financing?

The purchase price of the property was

?

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)

The cost of financing is

?

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)
Answer #1

Down Payment = D = $ 9127, Annual Payments = $ 768, Number of Payments = 9, Interest Rate = 9% compounded quarterly,

Equivalent Annual Interest Rate = [1+(0.09/4)]^(4)-1 = 0.09308 or 9.308 %

Property Price = D + Total Pesent Value of Annual Repayments = 9127 + 768 x (1/0.09308) x [1-{1/(1.09308)^(9)}] = 9127 + 4547.1931 = $ 13674.1931 ~ $ 13674.19

Cost of Financing = Interest Paid = Total Payments Made - Property Cost = 768 x 9 + 9127 - 13674.19 = $ 2364.8069 ~ $ 2364.81